# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v3

@@ 4.1. Experimental Setup
  • HumanEval+ (Liu et al., 2023): An extended version of the original HumanEval(Chen et al., 2021b) dataset, comprising 164 hand-crafted Python programming problems. The number of test cases has been increased by nearly 81×, enabling a more comprehensive evaluation of the functional correctness of generated code.
  • MBPP+ (Liu et al., 2023): A refined and optimized subset of the original MBPP(Austin et al., 2021) dataset, comprising 378 Python programming problems. MBPP+ reduces the number of problems compared to MBPP but increases the size of the unit tests by 35×, providing a much stricter evaluation of the functional correctness of LLM-generated code.
@@
Baselines. We compare AdaDec against two standard decoding strategies:
@@
  • AdapT (Zhu et al., 2023): An uncertainty-guided adaptive decoding strategy that adjusts the sampling temperature for tokens at the beginning of a line (high-uncertainty tokens) to increase randomness and help overcome difficult-to-predict points, while keeping low-uncertainty tokens more deterministic. For our experiments, we set the parameters as follows: $a=0.5$ for high-uncertainty tokens and $b=0.05$ for all other tokens.
@@ 4.2. RQ1: Effectiveness in Improving Functional Correctness
Table 2 presents the Pass@1 performance of three decoding strategies, Greedy, AdapT, and AdaDec, across seven LLMs evaluated on the HumanEval+, MBPP+ and DevEval datasets.
Table 2. Pass@1 comparison of Greedy, AdapT, and AdaDec on HumanEval+, MBPP+, and DevEval. Values in parentheses indicate the change relative to Greedy.
@@
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 51.83% 52.44%(+0.61%) 57.32%(+5.49%) 41.01% 41.27%(+0.26%) 49.47%(+8.46%) 13.45% 13.03%(-0.42%) 18.49%(+5.04%)
DS-6.7B 62.80% 63.41%(+0.61%) 65.24%(+2.44%) 58.47% 58.73%(+0.26%) 61.11%(+2.64%) 28.57% 28.57%(0.00%) 27.73%(-0.84%)
ST-3B 48.78% 51.83%(+3.05%) 50.61%(+1.83%) 38.89% 39.42%(+0.53%) 46.03%(+7.14%) 15.55% 13.03%(-2.52%) 22.69%(+7.14%)
QW-0.6B 17.07% 18.29%(+1.22%) 22.56%(+5.49%) 9.52% 11.11%(+1.59%) 30.42%(+20.90%) 2.94% 2.52%(-0.42%) 7.56%(+4.62%)
QW-1.7B 40.24% 42.07%(+1.83%) 44.51%(+4.27%) 37.57% 36.51%(-1.06%) 42.86%(+5.29%) 7.98% 7.56%(-0.42%) 11.76%(+3.78%)
QW-4B 53.05% 53.66%(+0.61%) 56.10%(+3.05%) 40.48% 41.01%(+0.53%) 51.32%(+10.84%) 9.24% 9.24%(0.0%) 16.81%(+7.57%)
QW-8B 56.71% 57.32%(+0.61%) 62.80%(+6.09%) 29.37% 29.63%(+0.26%) 47.09%(+17.72%) 4.62% 4.20%(-0.42%) 7.56%(+2.94%)
AVG. $\Delta$ +1.22% +4.09% +0.34% +10.43% -0.60% +4.32%
@@
Comparison with Greedy Search. Across all three benchmarks, AdaDec consistently outperforms Greedy decoding across most datasets and models. On HumanEval+, AdaDec improves pass@1 by an average of +4.09%, with gains as high as +6.09% for QW-8B and +5.49% for DS-1.3B. On MBPP+, the improvement is even more pronounced: AdaDec yields an average gain of +10.43%, with the most substantial boost observed for QW-0.6B (+20.90%) and QW-8B (+17.72%). On DevEval, AdaDec also yields meaningful improvements for most models (+4.32% on average), with ST-3B and QW-4B showing particularly strong gains (+7.14% and +7.57%, respectively). The only exception is DS-6.7B, where we observe a slight drop ($0.84\%$). This may be due to DevEval’s repository-level nature, where future logic are harder to capture with short-horizon lookahead, and DS-6.7B already performs strongly under Greedy decoding, leaving less room for improvement.
@@
Comparison with AdapT. When compared to AdapT, AdaDec again shows clear advantages. While AdapT offers marginal improvements over Greedy decoding (e.g., +1.22% on HumanEval+ and +0.34% on MBPP+), it sometimes underperforms (e.g., -0.60% on DevEval). In contrast, AdaDec consistently achieves larger gains, surpassing AdapT by wide margins across nearly all models and datasets. For example, on MBPP+, AdaDec exceeds AdapT by more than +10% on average, with the gap reaching +19.31% for QW-0.6B. On HumanEval+, AdaDec achieves substantially higher gains (average +4.09% vs. +1.22%), highlighting its stronger effectiveness in enhancing code correctness.
@@
Summary: Overall, the results indicate that AdaDec is highly effective in improving the functional correctness of generated code. Compared with Greedy search, it achieves consistent and often substantial improvements across most models and benchmarks. Compared with AdapT, AdaDec demonstrates more robust and scalable gains. These findings confirm the effectiveness of the proposed approach in enhancing decoding quality.
@@ 4.3. RQ2: Efficiency in Computational Cost
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval+, MBPP+ and DevEval datasets from the following perspectives. First, we measure the pause rate of AdaDec to assess the possible overhead introduced by its pause-then-rank mechanism. Second, we compare the prediction latency of AdaDec against Greedy decoding and AdapT method.
Pause Rate. As presented in Table 3, the pause rate ranges from 4.80% to 12.17%, with an average of 7.59%. This low activation frequency indicates that AdaDec concentrates computational effort on a small subset of high-uncertainty steps, effectively balancing local correction with overall efficiency. Across all datasets, the pause rate remains below 10% for all models except QW-1.7B, demonstrating that selective triggering successfully avoids the high cost associated with global lookahead.
@@
DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
HumanEval+ 6.51% 5.48% 4.80% 6.11% 11.17% 7.82% 5.98%
MBPP+ 8.90% 7.46% 6.52% 8.50% 12.17% 7.11% 8.33%
DevEval 7.41% 6.17% 6.59% 6.16% 11.47% 9.69% 5.04%
AVG. $\Delta$ 7.61% 6.37% 5.97% 6.92% 11.60% 8.21% 6.45%
Latency. Table 4 reports the average generation time per problem for Greedy, AdapT, and AdaDec across HumanEval+, MBPP+, and DevEval. Overall, AdaDec introduces moderate additional latency compared to Greedy, reflecting the computational overhead of the pause-and-rerank mechanism. On most models and datasets, the increase remains reasonable—for instance, +0.23s on HumanEval+ and -0.84s on MBPP+ on average—demonstrating that selective triggering effectively balances enhanced generation quality with efficiency.
AdapT exhibits mixed runtime behavior. In some cases, it is faster than Greedy (e.g., DS-1.3B on MBPP+), likely because higher-temperature sampling applied only at high-uncertainty tokens (typically line-start tokens) can lead to shorter, simpler continuations, reducing the total number of decoding steps. This effect is particularly noticeable on smaller models and in datasets like DevEval. In contrast, AdaDec performs targeted lookahead at every high-uncertainty step, which increases runtime, especially for larger models and more complex DevEval problems, but provides more reliable improvements in functional correctness.
Table 4. Comparison of the average generation time per problem (in seconds) by Greedy, AdapT, AdaDec on HumanEval+, MBPP+, DevEval (the brackets indicate the change ratio relative to Greedy)
@@
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 1.15 2.42 1.46 0.54 0.63 0.97 4.74 9.61 10.24
DS-6.7B 2.37 3.68 3.44 1.02 4.24 1.30 15.54 11.55 26.08
ST-3B 1.70 3.02 2.06 0.59 0.81 0.76 6.87 8.00 16.72
QW-0.6B 6.22 4.07 5.25 10.02 6.14 2.89 20.02 13.86 21.18
QW-1.7B 8.62 4.23 5.27 2.39 2.54 2.28 18.21 13.96 30.33
QW-4B 5.41 7.57 8.06 4.09 3.76 3.37 28.17 17.94 43.97
QW-8B 5.39 4.73 6.94 2.84 4.83 4.01 64.56 16.80 68.08
AVG. $\Delta$ -0.16 +0.23 +0.21 -0.84 -9.48 +8.36
Summary: AdaDec maintains high efficiency by selectively applying pause-and-rerank at a small fraction of high-uncertainty steps, with pause rates mostly below 10%. While this targeted lookahead adds moderate latency compared to Greedy, it remains manageable and effectively balances quality and efficiency. AdapT can sometimes reduce runtime via stochastic sampling, but AdaDec delivers more consistent improvements in functional correctness.
@@ 4.4. RQ3: Ablation Study
Predictive Effectiveness of Learned Entropy Thresholds. Table 5 summarizes the performance of a logistic regression classifier trained with learned entropy thresholds ($\tau^{\bm{LM}}$) across various language models. Overall, the classifier demonstrates strong effectiveness in predicting whether the top-1 token matches the ground truth. The highest performance is observed on DS-6.7B, achieving an accuracy of 91.05%, an F1 score of 0.9512, and an AUC of 0.9305. Other models, including DS-1.3B and ST-3B, also perform well, with F1 scores above 0.94 and AUC values exceeding 0.89. These results support the effectiveness of learned entropy thresholds as a reliable and lightweight signal for triggering pauses and reranking.
@@
QW-0.6B 0.8512 0.8774 0.9498 0.9122 0.8914 1.3213
QW-1.7B 0.8664 0.9003 0.9457 0.9225 0.8964 0.6153
QW-4B 0.8807 0.9166 0.9471 0.9316 0.9073 0.7134
QW-8B 0.8843 0.9158 0.9533 0.9342 0.9092 0.7060
AVG. $\Delta$ 0.8852 0.9117 0.9591 0.9348 0.9120 0.9393
Practical Effectiveness of Learned Entropy Thresholds. To further assess the impact of learned thresholds in practice, we compare the decoding performance of models using the learned threshold versus a fixed threshold. The fixed threshold is set to 1.2. Table 6 reports pass@1 results on HumanEval+. The experiments show that using the learned thresholds consistently improves performance compared to the fixed threshold. For DS-1.3B, DS-6.7B, ST-3B, and QW-0.6B, the learned thresholds are close to 1.2, resulting in modest gains. In contrast, for QW-1.7B, QW-4B, and QW-8B, the learned thresholds differ more from 1.2, yielding more noticeable improvements in decoding performance. This confirms that the learned thresholds not only perform well in predictive tasks but also translate into tangible gains during actual decoding.
@@
Strategy DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B AVG
Fixed 57.32% 65.24% 50.00% 21.34% 41.46% 51.22% 61.59% 49.74%
Learned 57.32% 65.24% 50.61% 22.56% 44.51% 56.10% 62.80% 51.31%
$\Delta$ 0.00% 0.00% +0.61% +1.22% +3.05% +4.88% +1.21% +1.57%
Summary: The learned entropy threshold $\tau^{\bm{LM}}$ is both theoretically sound and practically effective: it reliably distinguishes uncertain decoding steps and improves decoding performance when compared to fixed-threshold baselines.

### v4

@@ 3.2.3. Threshold Estimation
Training Cost Analysis. It is worth highlighting that this entire threshold-learning setup is computationally lightweight. Quantitatively, for most models, the process requires less than 1.5 GPU-hours, a cost that is dominated by the inference-only next-token prediction on BigCodeBench (Section 3.2.1). Crucially, this is a one-time, offline cost that involves no expensive parameter updates (e.g., retraining or fine-tuning), and the learned threshold $\tau^{\bm{LM}}$ is reusable for all subsequent inference tasks.
@@ 4.1. Experimental Setup
  • HumanEval+ (Liu et al., 2023): An extended version of the original HumanEval(Chen et al., 2021) dataset, comprising 164 hand-crafted Python programming problems. The number of test cases has been increased by a factor of nearly 81, enabling a more comprehensive evaluation of the functional correctness of generated code.
  • MBPP+ (Liu et al., 2023): A refined and optimized subset of the original MBPP(Austin et al., 2021) dataset, comprising 378 Python programming problems. MBPP+ reduces the number of problems compared to MBPP but increases the size of the unit tests by a factor of 35, providing a much stricter evaluation of the functional correctness of LLM-generated code.
@@
Baselines. We compare AdaDec against three decoding strategies:
@@
  • Beam Search (Freitag and Al-Onaizan, 2017): A classic decoding strategy that maintains the top-$k$ most probable sequences (beams) at each decoding step. To ensure a fair comparison with the lookahead mechanism in AdaDec, we set the beam size to $B=3$, matching the candidate pool size used in our method. This baseline helps distinguish the benefits of our uncertainty-guided reranking from those of simply exploring a wider search space.
  • AdapT (Zhu et al., 2023): An uncertainty-guided adaptive decoding strategy that adjusts the sampling temperature for tokens at the beginning of a line (high-uncertainty tokens) to increase randomness and help overcome difficult-to-predict points, while keeping low-uncertainty tokens more deterministic. For our experiments, we set the parameters as follows: $a=0.5$ for high-uncertainty tokens and $b=0.05$ for all other tokens. These values were determined through empirical tuning within the range suggested by the original study, selecting the configuration that yielded the best overall performance on our benchmarks.
@@ 4.2. RQ1: Effectiveness in Improving Functional Correctness
Table 2 presents the Pass@1 performance of four decoding strategies—Greedy, Beam Search ($B=3$), AdapT, and AdaDec —across nine LLMs evaluated on the HumanEval+, MBPP+, and DevEval datasets.
Table 2. Pass@1 (%) comparison of Greedy, Beam Search, AdapT, and AdaDec on HumanEval+, MBPP+, and DevEval. Values in parentheses indicate the change relative to Greedy.
@@
Greedy Beam AdapT AdaDec Greedy Beam AdapT AdaDec Greedy Beam AdapT AdaDec
DS-1.3B
62.20
62.20
(+0.00)
63.41
(+1.21)
65.24
(+3.04)
52.12
51.59
(-0.53)
52.65
(+0.53)
53.17
(+1.05)
14.29
18.91
(+4.62)
12.61
(-1.68)
18.49
(+4.20)
DS-6.7B
71.34
76.83
(+5.49)
72.56
(+1.22)
73.78
(+2.44)
65.61
65.87
(+0.26)
66.93
(+1.32)
67.72
(+2.11)
25.21
28.57
(+3.36)
26.47
(+1.26)
26.47
(+1.26)
ST-3B
48.78
50.61
(+1.83)
51.83
(+3.05)
50.61
(+1.83)
38.89
45.77
(+6.88)
39.42
(+0.53)
46.03
(+7.14)
15.55
13.03
(-2.52)
13.03
(-2.52)
22.69
(+7.14)
QW2.5-1.5B
64.02
68.90
(+4.88)
64.02
(+0.00)
67.07
(+3.05)
59.79
61.64
(+1.85)
60.32
(+0.53)
62.17
(+2.38)
10.08
11.34
(+1.26)
10.50
(+0.42)
14.71
(+4.63)
QW2.5-7B
84.15
83.54
(-0.61)
87.20
(+3.05)
87.20
(+3.05)
71.16
71.96
(+0.80)
72.22
(+1.06)
72.22
(+1.06)
18.49
14.71
(-3.78)
18.07
(-0.42)
23.11
(+4.62)
QW3-0.6B
17.07
24.39
(+7.32)
18.29
(+1.22)
22.56
(+5.49)
9.52
21.16
(+11.64)
11.11
(+1.59)
30.42
(+20.90)
2.94
5.88
(+2.94)
2.52
(-0.42)
7.56
(+4.62)
QW3-1.7B
40.24
41.46
(+1.22)
42.07
(+1.83)
44.51
(+4.27)
37.57
43.39
(+5.82)
36.51
(-1.06)
42.86
(+5.29)
7.98
9.24
(+1.26)
7.56
(-0.42)
11.76
(+3.78)
QW3-4B
53.05
57.32
(+4.27)
53.66
(+0.61)
56.10
(+3.05)
40.48
42.86
(+2.38)
41.01
(+0.53)
51.32
(+10.84)
9.24
13.45
(+4.21)
9.24
(+0.00)
16.81
(+7.57)
QW3-8B
56.71
59.76
(+3.05)
57.32
(+0.61)
62.80
(+6.09)
29.37
36.24
(+6.87)
29.63
(+0.26)
47.09
(+17.72)
4.62
6.30
(+1.68)
4.20
(-0.42)
7.56
(+2.94)
AVG. $\Delta$ +3.05 +1.42 +3.59 +4.00 +0.59 +7.61 +1.45 -0.47 +4.53
@@
Comparison with Greedy Search. Across all three benchmarks, AdaDec consistently outperforms Greedy decoding. On HumanEval+, AdaDec improves Pass@1 by an average of +3.59%, with gains as high as +6.09% for QW3-8B and +5.49% for QW3-0.6B. On MBPP+, the improvement is even more pronounced: AdaDec yields an average gain of +7.61%, with particularly significant performance boosts observed for QW3-0.6B (+20.90%) and QW3-8B (+17.72%). On DevEval, AdaDec demonstrates robust effectiveness, yielding meaningful improvements for all evaluated models (+4.53% on average). Particularly, QW3-4B and ST-3B show strong gains of +7.57% and +7.14%, respectively. Unlike the mixed results sometimes observed with other strategies, AdaDec achieves positive improvements across all models on DevEval, highlighting its ability to handle repository-level contexts where long-range logic consistency is crucial.
Comparison with Beam Search. A critical question is whether the gains of AdaDec stem merely from searching multiple candidate tokens or from our uncertainty-guided strategy. Comparing AdaDec with standard Beam Search (using the same beam width $B=3$), we observe that AdaDec consistently achieves superior performance. While Beam Search provides a baseline improvement over Greedy (averaging +3.05% on HumanEval+ and +4.00% on MBPP+), AdaDec surpasses it by further margins. The advantage of AdaDec is particularly evident on the more challenging MBPP+ and DevEval benchmarks. For instance, on DevEval, Beam Search yields a modest average gain of +1.45%, whereas AdaDec achieves +4.53%, nearly tripling the improvement. This indicates that blindly searching for top-$k$ continuations (as in Beam Search) is less effective than AdaDec’s approach of triggering lookahead only at high-uncertainty points, which allows for more targeted and semantically meaningful corrections.
@@
Comparison with AdapT. When compared to AdapT, AdaDec shows clear advantages in both stability and magnitude of improvement. While AdapT offers marginal improvements over Greedy decoding (e.g., +1.42% on HumanEval+ and +0.59% on MBPP+), it exhibits instability, leading to performance drops on DevEval for several models (average -0.47%). In contrast, AdaDec consistently achieves larger gains, surpassing AdapT by wide margins across nearly all models and datasets. For example, on MBPP+, AdaDec exceeds AdapT by approximately +7% on average. These results confirm that utilizing Shannon entropy to guide lookahead is a more reliable strategy than the heuristic-based temperature adjustment used in AdapT.
@@
Summary: Overall, the results indicate that AdaDec is highly effective in improving the functional correctness of generated code. It consistently outperforms Greedy decoding across diverse models and benchmarks. More importantly, it surpasses Beam Search, particularly on complex tasks (MBPP+ and DevEval), proving that uncertainty-guided intervention is superior to blind search. Compared with AdapT, AdaDec demonstrates significantly more robust and scalable gains.
@@ 4.3. RQ2: Efficiency in Computational Cost
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval+, MBPP+, and DevEval datasets. We analyze two key metrics: the pause rate and prediction latency.
Pause Rate. Table 3 presents the pause rates across different models and datasets. The results show that AdaDec triggers the reranking mechanism sparingly, with pause rates ranging from 0.85% to 12.17%, averaging approximately 6.75% across all configurations. This low activation frequency indicates that the entropy-based threshold successfully identifies a small subset of critical, high-uncertainty steps. Unlike global search strategies, AdaDec concentrates computational effort only where necessary. Notably, even for the largest models, the pause rate generally remains below 10%, demonstrating that our selective triggering mechanism scales efficiently without incurring the cost of constant lookahead.
@@
DS-1.3B DS-6.7B ST-3B QW2.5-1.5B QW2.5-7B QW3-0.6B QW3-1.7B QW3-4B QW3-8B
HumanEval+ 2.81% 2.33% 4.80% 1.48% 0.85% 6.11% 11.17% 7.82% 5.98%
MBPP+ 8.49% 7.46% 6.52% 6.24% 4.43% 8.50% 12.17% 7.11% 8.33%
DevEval 7.75% 6.29% 6.59% 6.95% 9.70% 6.16% 11.47% 9.69% 5.04%
AVG. $\Delta$ 6.35% 5.36% 5.97% 4.89% 4.99% 6.92% 11.60% 8.21% 6.45%
Latency. Table 4 reports the average generation time per problem for Greedy, Beam Search ($B=3$), AdapT, and AdaDec.
Comparison with Greedy. Overall, AdaDec introduces moderate additional latency compared to Greedy decoding (e.g., an average increase of +0.75s on HumanEval+ and +0.32s on MBPP+). This overhead is expected due to the lookahead computations. However, considering the substantial accuracy gains reported in RQ1, this trade-off is highly favorable.
Notably, for certain models (e.g., QW3-0.6B and QW3-1.7B on HumanEval+), AdaDec actually exhibits lower latency than Greedy decoding. This counter-intuitive result occurs because AdaDec prevents the model from drifting into erroneous, low-quality trajectories—such as repetitive loops or verbose, nonsensical outputs—that often plague Greedy decoding at high-uncertainty points. By correcting the generation path early via lookahead, AdaDec produces more concise and correct solutions, thereby reducing the total number of generated tokens and effectively offsetting the computational cost of reranking.
Comparison with Beam Search. A critical finding is that AdaDec is significantly more efficient than Beam Search, despite outperforming it in accuracy. As shown in Table 4, Beam Search introduces substantial latency overhead due to maintaining multiple active beams at every step. For instance, on the repository-level DevEval benchmark, Beam Search increases the average generation time by +24.62s compared to Greedy. In contrast, AdaDec increases it by only +7.47s. This efficiency stems from our selective pausing strategy: by triggering the lookahead mechanism only at high-uncertainty steps rather than globally, AdaDec avoids wasted computation on trivial tokens.
Comparison with AdapT. AdapT exhibits mixed runtime behavior. In simpler scenarios (e.g., MBPP+), it adds negligible overhead (+0.10s) or even reduces time slightly for specific models by promoting shorter sequences via sampling. However, AdaDec provides a more consistent balance: while slightly slower than AdapT, it delivers far more robust functional correctness (as shown in RQ1), justifying the marginal increase in runtime.
Table 4. Comparison of the average generation time per problem (in seconds) by Greedy, Beam Search, AdapT, AdaDec on HumanEval+, MBPP+, DevEval
@@
Greedy Beam AdapT AdaDec Greedy Beam AdapT AdaDec Greedy Beam AdapT AdaDec
DS-1.3B 2.94 7.01 2.72 4.56 1.46 1.91 1.47 1.68 4.03 14.59 4.24 7.66
DS-6.7B 9.55 25.60 13.59 11.87 3.21 7.51 3.13 4.81 20.56 53.52 23.72 32.91
ST-3B 1.71 3.45 2.72 2.06 0.60 1.01 0.84 0.77 6.92 34.40 8.33 14.90
QW2.5-1.5B 5.29 9.39 4.18 5.82 1.78 2.24 1.55 2.24 7.91 19.02 7.88 10.12
QW2.5-7B 12.24 38.22 13.40 14.38 3.94 10.12 3.90 6.02 12.12 26.51 13.71 22.19
QW3-0.6B 6.34 8.94 5.18 5.03 5.94 9.44 6.14 3.89 19.49 23.79 17.86 21.84
QW3-1.7B 8.23 9.98 6.03 5.29 2.13 3.17 2.54 2.28 22.21 46.98 20.56 29.53
QW3-4B 5.41 9.13 7.29 8.10 3.49 5.02 3.76 3.37 29.57 80.24 31.94 42.78
QW3-8B 5.62 11.76 5.63 6.99 3.88 5.19 4.03 4.21 61.14 106.48 63.53 69.23
AVG. $\Delta$ +7.35 +0.38 +0.75 +2.13 +0.10 +0.32 +24.62 +0.87 +7.47
Summary: AdaDec maintains high efficiency by selectively triggering lookahead at a low rate (average $\approx$ 6.75%), effectively filtering out low-utility computations. Crucially, compared to Beam Search, AdaDec incurs significantly lower additional overhead while achieving superior output quality across all benchmarks. This confirms that uncertainty-guided lookahead represents a far more efficient allocation of inference resources than blind search.
@@ 4.4. RQ3: Ablation Study
Predictive Effectiveness of Learned Entropy Thresholds. Table 5 summarizes the performance of the logistic regression classifier trained with learned entropy thresholds ($\tau^{\bm{LM}}$). Overall, the classifier demonstrates robust predictive capability across all models, achieving an average F1 score of 0.9351 and an AUC of 0.9116. The highest performance is observed on DS-6.7B, with an accuracy of 91.05% and an F1 score of 0.9512. Even for smaller models, the classifier remains reliable; for instance, ST-3B achieves an F1 score of 0.9475. These high metrics confirm that Shannon entropy, when calibrated via our learning process, serves as a high-fidelity signal for distinguishing between correct and incorrect token predictions.
@@
QW2.5-1.5B 0.8824 0.9055 0.9638 0.9337 0.9085 1.2232
QW2.5-7B 0.8889 0.9106 0.9679 0.9384 0.9119 1.1739
QW3-0.6B 0.8512 0.8774 0.9498 0.9122 0.8914 1.3213
QW3-1.7B 0.8664 0.9003 0.9457 0.9225 0.8964 0.6153
QW3-4B 0.8807 0.9166 0.9471 0.9316 0.9073 0.7134
QW3-8B 0.8843 0.9158 0.9533 0.9342 0.9092 0.7060
AVG. $\Delta$ 0.8853 0.9109 0.9606 0.9351 0.9116 0.9969
Practical Effectiveness of Learned Entropy Thresholds. To validate the necessity of learning the threshold rather than using a heuristic constant, we compare the decoding performance of models using the learned threshold versus a fixed threshold. The fixed threshold is set to 1.2. Table 6 reports the Pass@1 results on HumanEval+.
The experiments demonstrate that using the learned thresholds yields consistent improvements over the fixed threshold baseline across all evaluated models. The average improvement is +1.76%, with gains ranging from +0.61% to +4.88%. Specifically, while most models (e.g., DS-1.3B, QW2.5-7B, QW3-8B) show stable moderate gains between 1.2% and 1.8%, distinctively larger boosts are observed for QW3-1.7B (+3.05%) and QW3-4B (+4.88%). This indicates that a fixed threshold of 1.2 is often suboptimal. By calibrating the threshold to each model’s specific uncertainty distribution, AdaDec ensures robust performance gains, effectively unlocking the model’s potential regardless of its specific architecture or scale.
@@
Strategy DS-1.3B DS-6.7B ST-3B QW2.5-1.5B QW2.5-7B QW3-0.6B QW3-1.7B QW3-4B QW3-8B AVG
Fixed 63.41% 73.17% 50.00% 65.85% 85.98% 21.34% 41.46% 51.22% 61.59% 57.11%
Learned 65.24% 73.78% 50.61% 67.07% 87.20% 22.56% 44.51% 56.10% 62.80% 58.87%
$\Delta$ +1.83% +0.61% +0.61% +1.22% +1.22% +1.22% +3.05% +4.88% +1.21% +1.76%
Summary: The learned entropy threshold $\tau^{\bm{LM}}$ is both theoretically sound and practically effective. It reliably predicts token risks and consistently outperforms a fixed-threshold baseline across all models, delivering an average Pass@1 improvement of +1.76% and maximizing the effectiveness of the uncertainty-guided intervention.